# Hongyi Du1, Jiarui Yao1, Xiaocheng Yang1, Denghui Zhang1,3, Yunzhu Li4, Heng Ji1 1

### v1

@@ 6 Discussions and Future Directions
LM’s creativity for benchmarking. Our experiments with EscapeBench reveal key insights into the capabilities and limitations of current language models in creative reasoning and tool use. Even the most advanced model, EscapAgent, requires more hints and takes twice as many steps as the average human to solve tasks. The benchmark demonstrates that these advanced models struggle with implicit goal identification and creative problem-solving, often relying heavily on extensive hints. This underscores a significant gap in the evaluation of AI agents: while analytical and practical intelligence are well-assessed, creative intelligence remains largely underexplored. Addressing this gap may require fundamentally improving how LMs organize knowledge, linking objects not just by words but by their affordances—how their properties and functions relate to potential uses.
Human-AI Collaboration. Human-AI collaboration in escape room puzzles holds the potential to significantly enhance problem-solving efficiency and introduce a new paradigm of learning. Humans contribute intuitive insights and creative ideas that language models are unlikely to generate independently, while AI excels in systematic tasks such as data analysis and information aggregation. Furthermore, this collaboration leverages the strength of chain-of-thought reasoning, a process through which AI organizes and evaluates information in a logical, step-by-step manner. Integrating human creativity with AI’s structured problem-solving capabilities together in the future might not only foster innovative strategies but also enable deeper learning experiences, making the problem-solving process more dynamic and efficient.
Theoretical Foundations for AI Creativity. Understanding the cognitive mechanisms behind human creativity is essential for designing AI systems that emulate or surpass human creative processes. Human creativity is a multifaceted phenomenon involving the generation of novel and valuable ideas, problem-solving, and adaptation to complex situations Dainys (2024). Neurologically, it arises from the interplay between stochastic neuronal noise and structured, learned information, driven by spontaneous brain activity fluctuations Malach (2024). This contrasts with AI’s reliance on data patterns and algorithmic processes. Boden identifies three AI creativity mechanisms: combining familiar ideas in novel ways, exploring conceptual spaces, and enabling transformative innovations Boden (1998). Our research shows that even advanced models like GPT-4o struggle with implicit goal identification and creative problem-solving, often requiring extensive prompting. The EscapeAgent framework significantly reduces this dependency and improves task-solving efficiency, indicating that these modules effectively overcome barriers to creative reasoning. However, the results also highlight the importance of the core model’s capabilities, as larger models like GPT-4o benefit more from the framework than smaller models. This suggests further research is needed to address current models’ limitations in generating truly novel ideas. Interdisciplinary approaches integrating psychology, neuroscience, and model architecture can advance agent creativity further.
Multimodal Integration. Expanding the escape room environment to include multimodal data, such as visual and voice cues, offers a promising avenue for enhancing both the agent’s performance and the realism of the scenarios. While integrating vision-language models could enable agents to interpret visual clues more naturally, such an extension would also require robust visual understanding and reasoning capabilities to handle the complexity of tasks effectively. Additionally, incorporating multimodal interactions presents opportunities to study how agents synthesize information across modalities, such as correlating visual patterns with instructions or adapting strategies based on dynamic, multimodal feedback. Future work should explore how to seamlessly integrate and harmonize these diverse data types into the benchmark, pushing the boundaries of agents’ ability to process, reason about, and act on streams of information.
Step RL for Creative Reasoning. Introducing reinforcement learning into the EscapeBench task is expected to enhance the accuracy and efficiency of the agent’s exploration in long-chain tasks. This implies that under the guidance of rewards and penalties, the agent can explore in the correct direction more quickly. Compared with existing end-to-end reinforcement learning schemes, which rely on the final completion of the escape room task as the ultimate reward, introducing step rewards—providing immediate feedback for each step of the model’s operation—could potentially accelerate convergence and foster creative advancements in the model. Specifically, the discrete steps in EscapeBench can be organized either as a continuous, logically coherent progression or as strategic, abrupt jumps reflecting non-linear reasoning. Step Rewards can also further draw on the idea of task decomposition to be structured hierarchically. By constructing such a framework, Hierarchical Reinforcement Learning could manage super-long reasoning chains, decompose complex tasks into manageable subtasks, and enable a more fine-grained exploration and evaluation of AI’s creative capabilities.
@@ 7 Conclusion
In this work, we introduce EscapeBench, the first benchmark for advancing LM’s creativity. Our results highlight that current LMs still lag in creative reasoning. However, the EscapeAgent framework effectively lowers the barrier to creative problem-solving and implicit goal identification, resulting in significant performance improvements. While these advancements are promising, challenges remain in further enhancing the creative capabilities of the language model itself. Future work may focus on integrating multi-modal perception and developing new reinforcement learning algorithms to inspire even greater creativity. Our work serves as an important first step, offering a robust environment for experimentation. Looking ahead, we believe that creative intelligence, beyond just analytical and practical capabilities, will play a key role in shaping the frontier of AI.
@@ D.2 Resource Setting
D.3 More Study Results
@@
In Figure 13, we present the results of action trial counts for a specific item across four additional models. We observe that EscapeAgent achieves a higher success rate within 10 trials, even though it does not always succeed on the first attempt. Furthermore, for smaller models like Phi-3 and Ministral, EscapeAgent occasionally outperforms BaseAgent even in terms of one-trial success rates. This highlights how our framework effectively lowers the barriers for creative reasoning, even for less capable language models.

### v2

@@ 7 Discussions and Future Directions
Our experiments reveal that even the most advanced language models within EscapeAgent require more hints and twice as many steps as the average human, exposing limitations in creative reasoning and tool use. The benchmark highlights that while analytical and practical intelligence is well-assessed, creative intelligence remains a critical gap. Addressing this gap may require enhancing LMs to link knowledge and objects through affordances—their properties and functions—to foster creativity.
Theoretical Foundations for AI Creativity.
Human creativity, characterized by generating novel ideas and adapting to complexity, arises from the interplay of stochastic neuronal noise and structured, learned information Dainys (2024); Malach (2024). In contrast, AI creativity relies on trained data patterns and algorithms. Boden identifies three mechanisms driving AI creativity: combining familiar ideas, exploring conceptual spaces, and enabling transformative innovations Boden (1998). Integrating insights from psychology and neuroscience may further enhance AI’s creativity.
Human-AI Collaboration.
Human-AI collaboration in EscapeBench may promote a new problem-solving paradigm by merging human intuition with AI’s systematic reasoning. Humans bring unique insights and ideas that AI might not generate, while AI excels in tasks like information aggregation and logical organization. This synergy fosters innovative strategies, improves efficiency, and creates opportunities for deeper learning, offering a dynamic and enriched problem-solving experience that bridges human creativity with AI’s structured problem-solving capabilities.
@@ 8 Conclusion
In this work, we introduce EscapeBench, the first benchmark for advancing LM’s creativity. Our results show that while LMs still lag in creative reasoning, the EscapeAgent framework improves innovative problem-solving and implicit goal identification. Despite these advancements, enhancing the models’ intrinsic creativity remains a challenge. Future work could explore integrating multi-modal perception and new reinforcement learning algorithms to foster greater creativity. Our work serves as an important first step, offering a robust environment for experimentation. Looking ahead, we believe that creative intelligence, beyond just analytical and practical capabilities, will play a key role in shaping the frontier of AI.
@@
Ethical Statement
In this research, we consider the following ethical issues related to our benchmark and agent design:
$\bullet$ Fairness: We ensure that EscapeBench is designed to provide equal evaluation opportunities for all agents, regardless of their underlying model architectures or training methodologies. The tasks and scenarios are crafted to assess creativity and problem-solving abilities without bias, promoting fairness in the benchmarking process. Additionally, we aim to avoid overfitting to specific agent strategies, ensuring a more generalizable and inclusive evaluation framework for future AI advancements. While our environment is robust, we caution against potential misuse and strongly encourage its fair and responsible use.
$\bullet$ Transparency: Our work incorporates Chain-of-Thought reasoning in the BaseAgent framework to improve the transparency and interpretability of the agent’s decision-making process. This approach makes it easier to attribute the reasoning behind each agent’s action. Additionally, we will fully release the benchmarking code, EscapeAgent design, and data to promote transparency in our evaluation process, ensuring that the broader research community can benefit from and build upon our work.
Acknowledgment
This research is based upon work supported by U.S. DARPA ITM Program No. FA8650-23-C-7316 and DARPA ECOLE Program No. #HR00112390060. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ D.2 Resource Setting
Figure 12: Ablation of difficulty through Hints Used.
Figure 13: Ablation of difficulty through Total Steps.
D.3 Ablation Study
We perform an ablation study on the total steps and used hints in Figure 12 and Figure 13. Generally, harder game settings require more steps and hints for an agent to solve. Since our difficulty setting depends solely on the granularity and usefulness of descriptions and feedback (see Table 1), our results demonstrate that the way the environment is presented can impact difficulty, even when the core game logic remains unchanged.
Figure 14: A comparison of action trial times distribution for a specific item before success.
D.4 Further Analysis
We further analyze the valid actions agents attempt on different items in scenes before successfully operating them in Figure 14. While the BaseAgent exhibits a higher first-attempt success rate, EscapeAgent achieves greater effectiveness by making multiple attempts, leading to a significantly higher overall success rate within 10 trials. This difference can be attributed to EscapeAgent’s strategy of proposing multiple viable actions simultaneously. Although this increases the likelihood of eventually succeeding by trying more actions, it does not prioritize the most-likely-to-succeed action first. As a result, the BaseAgent appears more efficient on its initial trial, but its superficial approach, as highlighted in Table 4, limits its overall performance. The EscapeAgent’s design effectively addresses this limitation by leveraging a more exploratory approach, which proves advantageous in complex scenarios requiring creativity.
D.5 More Study Results
@@
In Figure 10, we present the results of action trial counts for a specific item across four additional models. We observe that EscapeAgent achieves a higher success rate within 10 trials, even though it does not always succeed on the first attempt. Furthermore, for smaller models like Phi-3 and Ministral, EscapeAgent occasionally outperforms BaseAgent even in terms of one-trial success rates. This highlights how our framework effectively lowers the barriers to creative reasoning, even for less capable language models.
@@
Appendix E Human Performance Details
To better understand the “Average Human” baseline we present in the main table, we conducted a post-experiment survey with all participants, including:
  • Prior offline Room Escape game experience? [Yes/No]
  • Number of times played offline? [Integer]
  • Prior online Room Escape experience (e.g., phone/PC)? [Yes/No]
  • Self-assessed skill level? [Skilled / Might Be / Not Really / Not at All]
  • Suggestions for improving EscapeBench
Results indicated that nearly all participants had prior experience, with an average of 5.6 offline sessions. However, only 20% identified as skilled players; 40% selected “Might Be” and another 40% “Not Really.” Despite varying self-assessments, game logs showed no significant performance difference across groups—likely due to the trial-and-error nature of the tasks.
Although a large-scale user study is beyond our current scope, these findings suggest that our reported human performance offers a fair approximation of average human reasoning under novel conditions.
Appendix F Scalability of Benchmark Creation
To ensure high-quality annotations, we employ human annotators supported by an automated checker and manual cross-validation. While we experimented with GPT-4o for annotation, it faced two major issues: (i) frequent omission of interactable objects, and (ii) inability to balance difficulty across scenarios. These limitations currently require human intervention to maintain benchmark quality.
Specifically, we explored using GPT-4o to generate item-based scenarios. While the model can produce coherent logical sequences, it lacks the creativity and complexity required by our benchmark. For example:
$\bullet$ Example 1: Use a magnifying glass to examine an ancient scroll, revealing hidden formulas.
$\bullet$ Example 2: Clean a rusty key with water, then scrape it with a coin to unlock a cabinet (partially creative).
Each AI-generated scenario yields about 12 key steps, whereas our benchmark includes over 1200 steps across 36 finely tiered scenarios. Therefore, human expertise remains essential for crafting creatively challenging and balanced content. While full automation falls short, parts of the process (e.g., validation, formatting) can be scaled. Future model improvements may increase automation potential, but human oversight will remain critical.
Appendix G More Detailed Discussions
We present a more detailed version of our results discussions and future research directions.
Theoretical Foundations for AI Creativity.
Understanding the cognitive mechanisms behind human creativity is essential for designing AI systems that emulate or surpass human creative processes. Human creativity is a multifaceted phenomenon involving the generation of novel and valuable ideas, problem-solving, and adaptation to complex situations Dainys (2024). Neurologically, it arises from the interplay between stochastic neuronal noise and structured, learned information, driven by spontaneous brain activity fluctuations Malach (2024). This contrasts with AI’s reliance on data patterns and algorithmic processes. Boden identifies three AI creativity mechanisms: combining familiar ideas in novel ways, exploring conceptual spaces, and enabling transformative innovations Boden (1998). Our research shows that even advanced models like GPT-4o struggle with implicit goal identification and creative problem-solving, often requiring extensive prompting. The EscapeAgent framework significantly reduces this dependency and improves task-solving efficiency, indicating that these modules effectively overcome barriers to creative reasoning. However, the results also highlight the importance of the core model’s capabilities, as larger models like GPT-4o benefit more from the framework than smaller models. This suggests further research is needed to address current models’ limitations in generating truly novel ideas. Interdisciplinary approaches integrating psychology, neuroscience, and model architecture can advance agent creativity further.
Multimodal Integration.
Expanding the escape room environment to include multimodal data, such as visual and voice cues, offers a promising avenue for enhancing both the agent’s performance and the realism of the scenarios. While integrating vision-language models could enable agents to interpret visual clues more naturally, such an extension would also require robust visual understanding and reasoning capabilities to handle the complexity of tasks effectively. Additionally, incorporating multimodal interactions presents opportunities to study how agents synthesize information across modalities, such as correlating visual patterns with instructions or adapting strategies based on dynamic, multimodal feedback. Future work should explore how to seamlessly integrate and harmonize these diverse data types into the benchmark, pushing the boundaries of agents’ ability to process, reason about, and act on streams of information.
Step RL for Creative Reasoning.
Introducing reinforcement learning into the EscapeBench task is expected to enhance the accuracy and efficiency of the agent’s exploration in long-chain tasks. This implies that under the guidance of rewards and penalties, the agent can explore in the correct direction more quickly. Compared with existing end-to-end reinforcement learning schemes, which rely on the final completion of the escape room task as the ultimate reward, introducing step rewards—providing immediate feedback for each step of the model’s operation—could potentially accelerate convergence and foster creative advancements in the model. Specifically, the discrete steps in EscapeBench can be organized either as a continuous, logically coherent progression or as strategic, abrupt jumps reflecting non-linear reasoning. Step Rewards can also further draw on the idea of task decomposition to be structured hierarchically. By constructing such a framework, Hierarchical Reinforcement Learning could manage super-long reasoning chains, decompose complex tasks into manageable subtasks, and enable a more fine-grained exploration and evaluation of AI’s creative capabilities.